OpenStudy (anonymous):

Will Medal and Fan Solve, using linear combination. 4x + y = 5 3x + y = 3 (0, –3) (0, 5) (2, –3) (–3, 7)

OpenStudy (anonymous):

@nincompoop ?

OpenStudy (anonymous):

@pooja195 ?

jimthompson5910 (jim_thompson5910):

You can subtract equation 2 from equation 1 to get 4x + y = 5 3x + y = 3 ---------- 1x + 0y = 2 solve for x and use this to find y

OpenStudy (anonymous):

So would this be B?

jimthompson5910 (jim_thompson5910):

when you isolated x, what is x equal to?

OpenStudy (anonymous):

I'm not sure.

jimthompson5910 (jim_thompson5910):

you are solving 1x + 0y = 2 for x

jimthompson5910 (jim_thompson5910):

0y is the same as 0 since 0 times anything is 0

jimthompson5910 (jim_thompson5910):

1x = x because 1 times anything is that same number

OpenStudy (anonymous):

Oh, so it's C?

OpenStudy (anonymous):

@jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

when you solved for x, you got x = ???

OpenStudy (anonymous):

I don't really understand.

jimthompson5910 (jim_thompson5910):

then how did you get to the answer of C? just a random guess?

OpenStudy (anonymous):

No I used the numbers from the answer C and put them into the X and Y spots

jimthompson5910 (jim_thompson5910):

that works, you can plug each possible answer choice into the original equations

OpenStudy (anonymous):

Thanks for the help :)

jimthompson5910 (jim_thompson5910):

you should get 2 true equations, so yes, C is the answer